A 25.00 mL sample of 0.685 M cobalt (II) phosphite (X) a. has a molar mass of 105.9064 g/mol b. has a molar mass of 105.906 g/mol 4. contains 1.8*10 g CX) contains 6.280 g contains 6.28 g CX) contains 6.3 g contains 5.732 g CX) contains. contains 5.l g contains 4.380 g CX) contains 4.38 g( contains 4.4 g contains 1.814 g CX) contains 1.81 g CX) contains 1.8 g (X) contains 2 PO4 ions contains 6 PO4 ions contains 1.37 PO, ions contains 4.11 PO4 ions c. has a molar mass of 105.91 g/mol d. has a molar mass of 105.9 g/mol e. has a molar mass of 106 g/mol f. has a molar mass of 255.7716 g/mol g. has a molar mass of 255.772 g/mol h. has a molar mass of 255.77 g/mol i. has a molar mass of 255.8 g/mol mm. nn. oo. pp. qq. j. has a molar mass of 256 g/mol k. has a molar mass of 334.7436 g/mol l. has a molar mass of 334.744 g/mol m. has a molar mass of 334.74 g/mol n. has a molar mass of 334.7 g/mol o. has a molar mass of 335 g/mol p. has a molar mass of 366.7424 g/mol q. has a molar mass of 366.742 g/mol r. has a molar mass of 366.74 g/mol uu. Vv ww. contains 3.43*102 PO4 ions xx. contains 1.03*101 PO4 ions yy. contains 2.06*10 PO ions zz. contains 6.19 1022 PO ions aaa. contains 3 Co ions bbb. contains 2 Co ions s. has a molar mass of 366.7 g/mol t. has a molar mass of 367 g/mol contains 6.280*10 g X) contains 6.28*10 g (X) contains 6.3*10 g (X) contains 5.732*103 g (X) contains 5.73 103 contains 5.7*10 g contains 4.380*103 g (X) V. w. x. contains 5.1410 Co ions g (X) ddd. contains 3.43*10 Co ions eee. contains 3.09*102 Co ions fff. contains 2.06*10 Co ions ggg. contains 8.53*1026 Co ions hhh. contains 5.69*1026 Co3 ions aa. bb. contains 4.38 10 g (X) cc. contains 4.4*10 g x) dd. ee. contains 1.814*10 g (X) contains 1.81*10 g (X)
Answer #1

Given complex concentration = 0.685 M

number of moles of complex = 0.685 * 25/1000 = 1.7125 * 10^-2 moles

name of the complex indicate it contains Co+2 and Phophite PO4 3- ions, so the formula of the complex will be Co3(PO4)2

molar mass of Co3(PO4)2 = 3*58.933 + 2*(30.9737 + (4*15.99)) = 366.7 g/mol

1 mole of Co3(PO4)2 contains 3 moles of Co+2 ions

1.7125 * 10^-2 moles of Co3(PO3)2 contains 3 * 1.7125 * 10^-2 moles of Co+2 ions

                                          = 5.1375 * 10^-2 moles of Co+2

mass of CO2+ ions = 5.1375 * 10^-2 * 58.933 = 3.03 g

number of Co+2 ions = 5.1375 * 10^-2 * 6.023 * 10^23 = 3.09 * 10^22 ions of Co+2

1 mole of Co3(PO4)2 contains 2 moles of PO4 3- ions

1.7125 * 10^-2 moles of Co3(PO3)2 contains 2 * 1.7125 * 10^-2 moles of PO4 3- ions

                                          = 3.425 * 10^-2 moles of PO4 3- ions

number of PO4 3- ions = 3.425 * 10^-2 * 6.023 * 10^23 = 2.06 * 10^22 ions of PO4 3-

mass of PO4 3- = 3.425 * 10^-2*(30.9737 + (4*15.99)) = 3.25 g